OpenStudy (anonymous):

write the series summation notation. 2+5+8+11+ . . .

OpenStudy (praxer):

$$\huge \sum_{k=0}^{n-1} 2+3k $$ as $$ a_n =a+(n-1)d $$ here n goes from 1 to n so, (n-1) will go from 0 to n-1 replacing n-1 = k we get, the formual to be 2+3k where k goes from 0 to n-1